OpenStudy (samigupta8):

We have two capacitor connected in series of same capacitance C.When the dielectric of dielectric constant K is inserted between the plates of one of the capacitor ,the total charge flowing through the battery is

OpenStudy (samigupta8):

I got the magnitude bt the direction part is not matching with the ans given in ans key.. Can anyone pls help?? My direction is from battery's positive terminal to -ve terminal bt they have given just the opposite of that

OpenStudy (samigupta8):

@michele_laino

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (samigupta8):

@vincent-lyon.fr

OpenStudy (samigupta8):

@baru

OpenStudy (samigupta8):

@priyar

OpenStudy (priyar):

Yes..i think you are right.. coz if we use the same battery then some ext. source(battery) is supplying the extra charge..

OpenStudy (samigupta8):

I just need the confirmation ...of the same ...

OpenStudy (samigupta8):

@hartnn

ganeshie8 (ganeshie8):

Inserting dielectric reduces the electric field and thus increases the capacitance by a factor of \(K\) : \[C_{new} = C*K\]

OpenStudy (samigupta8):

Sir BT we know that final charge on the capacitor is kCE/k+1 which is greater than C/2.. So isn't the battery supplying the charge to the capacitor

ganeshie8 (ganeshie8):

\[C_{effective} = \dfrac{C(CK)}{C+CK} = \dfrac{CK}{1+K} \]

ganeshie8 (ganeshie8):

Right, that is greater than \(\dfrac{C}{2}\) if \(K\gt 1\)

OpenStudy (samigupta8):

So that means the capacitor is getting more charged or better we can say that battery has given it a charge this much k-1/2(k+1) CE

ganeshie8 (ganeshie8):

You're right and the textbook is wrong

OpenStudy (samigupta8):

Battery's positive terminal to -ve terminal Ryt?

ganeshie8 (ganeshie8):

Since the capacitance has increased, the charge on each of the plates of the capacitor also increases. The energy required for this charge movement is supplied by the battery. So yeah, the energy flows from battery to the capacitors

ganeshie8 (ganeshie8):

depends on what charge carriers you're considering

ganeshie8 (ganeshie8):

In this case, the battery is charging the capacitors. So, electrons move from negative terminal of battery to the positive terminal, external to the battery.

ganeshie8 (ganeshie8):

remember, the charge carriers in metals are electrons

OpenStudy (samigupta8):

Well! That's not mentioned actually.. About what charge are they demanding the ans .. Yep...now I got that ..I were considering positive charge carriers

ganeshie8 (ganeshie8):

then you're right "negative charge moving from A to B" is equivalent to "positive charge moving from B to A" for many calculations

OpenStudy (samigupta8):

Tysm sir...

ganeshie8 (ganeshie8):

Np

OpenStudy (irishboy123):

mmmm!!!! putting the dielectric in means that you increase the capacitance of the circuit so more charge should flow, **by Benjamin Franklin definition**, from the +ve terminal. that is conventional current. dunno, but just thought it might be worth saying :p

OpenStudy (priyar):

yeah the direction of current is what you have mentioned.. the flow of charges is just the reverse..!

OpenStudy (samigupta8):

@irishboy123 exactly more +ve charge will definitely flow from positive terminal to -ve terminal.. Bt this means that electrons are going to move from -ve terminal to +ve terminal... And charge carriers as mentioned by @ganeshie8 are electrons so we can say that charge flows from negative terminal to +ve terminal

OpenStudy (irishboy123):

well that depends and that was kinda my point

OpenStudy (samigupta8):

So as per you the charge flows from +ve to -ve side